Rose Co. had a net loss of $150,000 in 2015 when the selling price per unit was $20, the variable costs per unit were $15, and the fixed costs were $600,000. Management expects per unit data and total fixed costs to be the same in 2016. Management has set a goal of earning net income of $75,000 in 2016.

Required:

(a)   Compute the units sold in 2015.

(b)   Compute the number of units that would have to be sold in 2016 to reach management's desired net income level.

(c)   Assume that Rose Co. sells the same number of units in 2016 as it did in 2015. What would the selling price have to be in order to reach the target net income? Use the mathematical equation.
Answer #1

Part 1

Net profit (loss) = (20*X)-(15*X)-600000

-150000=20X-15X-600000

450000=5X

X= 90000

Units sold = 90000

Part 2

Desired sales = (fixed expense + desired earning) / contribution margin per unit = (600000+75000)/(20-15) = 135000 units

Contribution margin = sales price per unit– variable expense per unit

Part 3

Desired sales = (fixed expense + desired earning) / contribution margin per unit = (600000+75000)/(20-15)

90000 =(600000+75000)/(X-15)

90000 = 675000 / (X-15)

(X-15) = 675000/90000

X-15 = 7.5

X = 7.5 +15

X = $22.50

selling price = $22.50 per unit